OpenStudy (anonymous):

solve the trig equation: secx+tanx=1

OpenStudy (anonymous):

please help

OpenStudy (anonymous):

help

OpenStudy (anonymous):

sec x = 1 / cos x tan x = sin x / cos x So we have \[\frac{ 1 }{ \cos x } + \frac{ \sin x }{ \cos x } = 1\] cos x = 1 + sin x Can you proceed from here?